Find the area under the normal curve in each of the following cases.

  1. between z = 0 and z = 1.63
  2. between z= 0 and z = 1.78
  3. between z = -2.46 and z = 1.55
  4. to the right of z = 2.35
  5. to right of z = -1.31
  6. to the left of z = 0.35
  7. to the left of z - 1.85
  8. to the right of z = -0.95
1
Expert's answer
2021-04-27T08:51:01-0400

1. P(0<Z<1.63) = P(Z<1.63) – P(Z<0)

= 0.9487 – 0.5

= 0.4484

2. P(0<Z<1.78) = P(Z<1.78) – P(Z<0)

= 0.9624 – 0.5

= 0.4624

3. P(-2.46<Z<1.55) = P(Z<1.55) – P(Z< -2.46)

= 0.9394 – 0.0069

= 0.9325

4. P(Z>2.35) = 1 – P(Z<2.35)

= 1 – 0.9906

= 0.0094

5. P(Z> -1.31) = 1 – P(Z < -1.31)

= 1 – 0.0951

= 0.9049

6. P(Z<0.35) = 0.6368

7. P(Z< -1.85) = 0.0321

8. P(Z> -0.95) = 1 – P(Z< -0.95)

= 1 – 0.1710

= 0.8290